OpenStudy (anonymous):

3|3x-5|>=66 i have...

OpenStudy (anonymous):

\[[9,\infty)\cup(-\infty,-\frac{ 17 }{ 3 }]\]

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

yes it is correct. consider when 3x-5 is positive, |3x-5|>=22 and when 3x-5 is negative, it has to be less than -22 so that the absolute value is bigger than 22.

OpenStudy (anonymous):

thank you, sorry I couldn't get to you earlier the site was having difficulties.

OpenStudy (anonymous):

lol i know...i had hard time getting in this page...it says loading all the time :D